write the equation of the line through (0,5) perpendicular to y= -2x+1
Answer by stanbon(75887)

write the equation of the line through (0,5) perpendicular to y= -2x+1
--
The given slope is -2
---
The slope you want is 1/2
--------------------------------
The form is y = mx + b
You know m = 1/2 and you know y = 5 when x = 0.
So solve for "b":
----
5 = (1/2)*0 + b
b = 5
------------------------
Equation:
y = (1/2)x + 5
===================
Cheers,
Stan H.